OpenStudy (anonymous):

A square field has an area of 479ft power 2. What is the approximate length of a side of the field? Give your answer to the nearest foot. Explain your response. (Essay Answer)

OpenStudy (anonymous):

Area=side^2 479=side^2 \[\sqrt{479}=side\] Does that help?

OpenStudy (anonymous):

Oh...ya thanks!!